# Rethinking Weight-Averaged Model-merging

### v3

@@ Abstract
Model-merging has emerged as a powerful approach in deep learning, capable of enhancing model performance without any training. However, the underlying mechanisms that explain its effectiveness remain largely unexplored. In this paper, we investigate this technique from three novel perspectives to empirically provide deeper insights into why and how weight-averaged model-merging works: (1) we examine the intrinsic patterns captured by the learning of the model weights, through the visualizations of their patterns on several datasets, showing that these weights often encode structured and interpretable patterns and that is the essential why model-merging can work; (2) we mathematically and empirically investigate model ensemble merging strategies based on averaging on weights versus averaging on features, providing detailed analyses across diverse architectures and datasets; and (3) we explore the impact on model-merging prediction stability in terms of changing the parameter magnitude, revealing insights into the way of weight averaging works as regularization by showing the robustness across different parameter scales. Our findings shed light on the “black box” of weight-averaged model-merging, offering valuable insights and practical recommendations that advance the model-merging process. The code is available at https://github.com/billhhh/Rethink-Merge.
@@ 1 Introduction
Model-merging can combine multiple independently trained models into a single, effective model without introducing extra inference computations. This family of approaches has been applied in contexts such as natural language processing [38], computer vision [16], and their sub-areas (e.g. federated learning [32], distillation [13], adversarial learning [44, 2] and large language model [22, 45]), where the ability to merge models effectively can lead to largely improved models’ performance, efficiency, and flexibility in deployment. Despite the practical importance of model-merging, there has been limited research on the investigation of the underlying factors that make it effective.
This research gap motivates the need to innovatively explore the underlying principles of why and how the family of weight-averaged model-merging methods can work. In this paper, we bridge this gap by providing a comprehensive analysis of weight-averaged model-merging from the following three key perspectives:
  • An examination of the patterns contained in the model weights through the visualization of the patterns present in both linear layers as classifiers or inside deep learning models used on multiple datasets. Results reveal that clear and structured patterns emerge within the weights, suggesting that weight averaging can be seen as a linear vector combination, which contributes to the success of merging models via weight operations;
  • An investigation of averaging on weights versus averaging on features strategies allows us to evaluate the relative effectiveness and limitations of model-merging and traditional model ensembles across multiple architectures and datasets, uncovering distinct benefits and behaviors;
  • An exploration of the robustness of model-merging predictions with respect to different weight magnitudes/variances, which provides insights into the model-merging works as a type of model regularization.
Based on our viewpoints and findings, we conclude with useful proposals to provide a good reference for the community.
@@ 2 Related Work
Figure 1: Average of all images for each class (a) and the linear classifier visualization for each class on Cifar100 [14] dataset (b). We train a single linear classifier for all classes then wrap it back to the size of input images. The detailed experimental analysis is in Sec. 6.1.
@@
In recent years, model-merging [28, 23, 3, 21, 25, 37] has attracted growing interest. This approach has led to a range of methods focused on merging weights from independently trained models to enhance model performance or generalization. For example, Wortsman et al. [35] proposed an approach of averaging, or greedily averaging, the parameters from various models to yield “Model Soups” that show more accurate performance than the original models used to produce the “soup” but without introducing more computations into inference. Further, Yadav et al. proposed Ties-Merging [36] based on parameter averaging by considering conflicts between the model weights, allowing for smooth parameter merging across models. Some other approaches [1, 6, 12] address the problem from a different perspective, focusing on taking weight merging operations via finding a good way to interpolate on the loss basins. Zipit [27] takes one step ahead from Git Re-Basin [1] to align weights from inter-model and intra-model based on the alignment of features between merging models. Yu et al. [41] proposed a DARE model to combine models trained on domain-specific data to create a domain-agnostic model-merging that generalizes well across diverse distributions. This approach ensures that domain knowledge is preserved while enhancing the model’s ability to generalize across different tasks. Beyond model-merging, other techniques, such as mode connectivity path ensembles [8], analyze pathways in the parameter space between models. By identifying optimal modes along these paths, these models help improve the performance after merging. Fisher Merging [19] selects parameters that approximately maximize the joint likelihood of the models’ parameter posteriors.
Despite these advances, research into the underlying mechanisms that make model-merging effective remains limited. This gap in understanding serves as the motivation for our paper. To focus on a specific scope, we aim to offer a thorough analysis from mathematical and empirical perspectives of weight-averaged model-merging different from the study in [35], which serves as the foundation work for the area.
@@ 3 The Patterns Contained in Model Weights
In this section, we will explain the reason why does the weight-averaged model-merging work via the analysis of their weights from the perspective of template matching (an example is shown Fig. 1 and Fig. 2 for a linear model; for nonlinear models, visualizations can be found in the supplementary materials). Specifically, we explain the reason that the weights can be viewed as templates learned from the data and build connections with an inner product which is a widely explored operation in neural networks. By treating the weights as templates, it is understandable that operations in weight space as efforts to manipulate feature representations.
@@ 3.1 Connections Between Inner Product and Template Matching
The weights of the model can be interpreted as templates or prototypes for the respective class activations and it has a deep connection with machine learning. A couple of decades ago, Adaboost [7] contained heuristically designed patterns such as HAAR features [31], which are simple hand-crafted filters that capture edges / lines / other low-level image features through weighted sums and differences in pixel intensities. During training, Adaboost iteratively adjusts the weights of weak classifiers to improve their performance on misclassified examples.
In a linear classification, once the weights are learned, we wrap it back to have the same size with the input images. Take Cifar100 for instance, each image within the dataset is 32$\times$32$\times$3, if we flatten it and it will become a 3072$\times$1 vector; the weight matrix size is 100$\times$3072 without considering bias; in this case, each row of the weight matrix corresponds to each class to be classified and thus each row vector (1$\times$3072) can be wrapped back to 32$\times$32$\times$3 with the same size of the original images. The learned weight matrix serves as simplified templates that represent each class within the data, as shown in Fig. 6.1.
A key mechanism enabling these templates to operate effectively is the use of inner products to apply them to the input data. When an image is presented to the model, its activation is determined by the inner product between the image features and per class weight vector. For each class, which weight vector yields the highest activation is selected as the predicted class of that vector, indicating that the “template” aligns (or similar to) the input image most closely. This phenomenon also holds for deeper networks, but differently from the linear classifier, weight templates in each layer are learned to detect different features that contribute the most towards the objectives (e.g. in the lower layer, low level features, e.g. edge / line features, are tend to be learned). The averaging operation in the weight space represents as Fig. 2, which can be seen as the Mixup [42] operation in the weight space. Therefore, the linear combinations in the weight space has a similar effect as in the feature space. This also explains why recent research [4, 43, 11, 46] applies “model arithmetic” on multiple models in the weight space can achieve the effect of feature arithmetic as Word2vec [20].
@@ 3.2 Formal Representation from the Perspective of Template Matching
where $\|\mathbf{w}\|$ and $\|\mathbf{x}\|$ are the norms of $\mathbf{w}$ and $\mathbf{x}$, and $\theta$ is the angle between the two vectors. When $\theta=0$, then $\cos\theta=1$, which means that we have $\mathbf{w}\cdot\mathbf{x}=\|\mathbf{w}\|\|\mathbf{x}\|$, resulting in a weight matrix $\mathbf{w}$ that perfectly matches the input $\mathbf{x}$, as the inner product in (1) is maximized when the template $\mathbf{w}$ is perfectly aligned with the input $\mathbf{x}$. On the other hand, when $\theta\neq 0$, then $\cos\theta<1$, which means that the template and image features become misaligned. This mechanism also applies for the layers within deep learning models. The visualization of deep models can be found in the supplementary materials.
@@
Figure 3: The 10 model-merging/ensemble accuracy comparison between different architectures on Cifar100. Comparison between averaging on Weights (Uniform Soups, Greedy Soups), averaging on features (Ens Logits, Ens Features, Greedy Ens Logits, Greedy Ens Features) and original model performance (Performance Average) on Cifar100 using many different deep learning architectures, including CNNs and ViTs. The detailed experimental analysis is in Sec. 6.2.
Averaging on weights (model-merging) and Averaging on features (model ensembling) are two different ways to aggregate knowledge from multiple models without retraining. These methods have distinct properties and lead to different testing behaviors. In this section, we explore the differences between these two approaches and analyze their effects on the model effectiveness and efficiency, one example experiment is shown in Fig. 3.
@@ 4.2 Non-linear Model Scenarios
In such a case, the equivalency between $\overline{\mathbf{h}}$ and $\overline{\mathbf{h}}^{\prime}$ seen in the linear models may not hold for non-linear models. Take ReLU for instance, if elements of $\mathbf{W}_{1}\mathbf{x}$ and $\mathbf{W}_{2}\mathbf{x}$ exhibit totally opposite activations (same absolute values but with different signs), where their average will tend towards zero, leading to large loss of information after the non-linear activation. Therefore, averaging on weights tends to smooth out the contributions from $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, which can be beneficial in certain scenarios by reducing variance or providing a regularizing effect. However, this smoothing may also lead to a loss while extracting fine-grained details. In contrast, averaging on features allows each weight matrix to independently influence the final output after activation. However, averaging on features (model ensemble) increases the computational costs linearly with the number of ensembled models, while averaging on weights (model-merging) does not suffer. The FLOPs of models considered in the paper can be found in the supplementary materials.
@@ 5 The Model Predictions and Weight Magnitudes
Figure 4: Scatter plot comparing the classification accuracies of model merge (uniform soup) and model ensemble (logits ensemble) over different weights magnitudes with VGG19 model on Cifar100. The values on the X and Y axis represent the accuracies. The experimental analysis is in Sec. 6.3.
In this section, we will examine how the weight-merging behaves against magnifying the candidate model weights by different magnitude factors. This helps us understand how robust model-merging works.
@@ 5.1 Analysis of Weight Magnitude After Averaging
We start from analyzing the effect of averaging the weights of two independent matrices, denoted as $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, on their respective magnitudes. We define the maximum entry of both matrices as $M=\max(\mathbf{W}_{1},\mathbf{W}_{2})=\max_{i,j}(|(\mathbf{W}_{1})_{i,j}|,|(\mathbf{W}_{2})_{i,j}|)$. When we compute the averaged weights, we obtain:
@@
Mathematical Justification. Using the triangle inequality [30], we establish:
@@
Thus, it follows that $\|\overline{\mathbf{W}}\|\leq M$ combining with Eq. 10. It confirms that averaging the weights results in a new weight whose entry’s upper bound is equal to the maximum entry of the original weights. Thus, the operation of averaging the weights $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ typically results in a reduction of the maximum weight magnitude. This conclusion also extends to complex deep neural networks.
@@ 5.3 How Weight Magnitudes and Variance Effects Model Outputs
We have the following property 1 and theorem 1 to build the connections between model weights’ magnitudes/variances and outputs’ magnitudes/variances, following  [34]:
@@
where $M_{l}$ denotes the maximum number of layers, and $L$ represents the Lipschitz constant. This inequality shows that the output variance grows with the depth of the network, as well as impacted by the variances of the weights / bias.
@@
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 3: As the aforementioned analyses and experimental study in Sec. 6.3, model-merging is more robust to the changes in weights’ magnitudes/variances than model ensemble, by bringing down the maximum magnitudes/variances of individual models before merging. This is because, generally, the small magnitudes of weights enhances robustness by minimizing the network’s sensitivity towards input variations, but somehow reduces the expressiveness from ensembling all models. From this perspective, model-merging acts similarly as regularization. This balance between robustness and expressiveness is crucial and should be carefully considered based on the specific task requirements. For certain tasks where noise resilience is more valuable than fine-grained feature sensitivity, lower weight magnitudes can improve performance.
@@ 6.1 The Patterns Contained in Model Weights
Figure 6: Average of all images for each class and linear classifiers visualization for each class in Tiny ImageNet [15] dataset.
As shown in Fig. 1(a,b), the visualization of the linear classifier on Cifar100 explicitly shows the patterns per classifier learned, which are strongly correlated with the average of images per class. The patterns learned per class acts as a fused representative image. Interestingly, if we merge the weights of the first 5 classes with the last 5 classes in Fig. 1, it gives us the results of Fig. 2. As mentioned, the weight-averaged model-merging acts as a Mixup operation, which is a linear combination of different matrices, in the weight space. For instance, when the “bicycle” weight matrix is combined with the “bottle” weight matrix, we can see the wheels of bikes combined with a vertical bottle. We can also observe that, after merging, each element within the combined template noticeably fades. This phenomenon can also explain the limitation of expressiveness of the merged model as described in Sec. 4.
Fig. 6 visualizes the image average per class and its corresponding weight for Tiny ImageNet classes. For instance, “goldfish” shows a reddish center in the averaged images and also shows a reddish center in the classifier weights; similarly, we can observe the eyes and nose of “koala”; for pretzel, we can see the shape of the cracker from its weight visualization. However, they are not as clear as for Cifar100. This may be due to the relatively large dataset size, as the average of all images in a class itself is also vaguer. This compromised clarity can also be caused by the limited expressiveness of single linear classifiers and there are more complicated patterns present in Tiny ImageNet than Cifar100. Besides the linear classifier visualization, we also conduct common deep learning models with nonlinear weights visualization by viewing the first layer of convolutional kernels (7 $\times$ 7 in size) from the ResNet50 model [9] in Fig. 7 (VGG model weight visualization can be found in the supplementary materials). We can see the patterns of texture and edges for the basic level geometric feature detection is similar to the Haar features described in Sec. 3.1.

### v4

@@ Abstract
Model-merging has emerged as a powerful approach in deep learning, capable of enhancing model performance without any training. However, the underlying mechanisms that explain its effectiveness remain largely unexplored. In this paper, we investigate this technique from three novel perspectives to empirically provide deeper insights into why and how weight-averaged model-merging Wortsman et al. (2022) works: (1) we examine the intrinsic patterns captured by the learning of the model weights, and we are the first to connect that these weights encode structured with why weight-averaged model merging can work; (2) we investigate averaging on weights versus averaging on features, providing analyses from the view of diverse architecture comparisons on multiple datasets; and (3) we explore the impact on model-merging prediction stability in terms of changing the parameter magnitude, revealing insights into the way of weight averaging works as regularization by showing the robustness across different parameter scales. The code is available at https://github.com/billhhh/Rethink-Merge.
@@ 1 Introduction
Model-merging can combine multiple independently trained models into a single, effective model without introducing extra inference computations. This family of approaches has been applied in contexts such as natural language processing Yang et al. (2024), computer vision Li et al. (2023), and their sub-areas (e.g. federated learning Wang et al. (2020), distillation Khanuja et al. (2021), adversarial learning Zhang et al. (2024); Croce et al. (2023) and in large language models Rame et al. (2024); Zhou et al. (2024)), where the ability to merge models effectively can lead to largely improved models’ performance, efficiency, and flexibility in deployment. Despite the practical importance of model-merging, there has been limited research on the investigation of the underlying factors that make it effective.
This research gap motivates the need to explore the underlying principles of why and how the family of weight-averaged model-merging methods can work. In this paper, we bridge this gap by providing a comprehensive analysis of weight-averaged model-merging from the following three key perspectives:
  • We are the first to connect the patterns contained in the model weights to model-merging. It is conducted through the visualization of the patterns present in both linear classifiers or inside deep models on multiple datasets. Results reveal that clear and structured patterns emerge within the weights, suggesting that weight averaging can be viewed as a linear combination of two weight vectors;
  • We present a structured comparison between weight averaging and feature averaging strategies, focusing on their implications for model merging and ensembling. Our analysis across diverse models and datasets reveals consistent behavioral differences, offering a more nuanced understanding of model merging and ensembling behavior with different architectures;
  • We systematically analyze how the robustness of model-merging predictions is affected by differences in weight magnitudes and variances. Our findings reveal that model merging implicitly acts as a form of regularization, favoring solutions with smoother or more stable parameter configurations. This connection provides a theoretical perspective on when and why model merging improves generalization.
The intention of the paper: Instead of providing a thorough evaluation of existing works, e.g. Ainsworth et al. (2022); Yadav et al. (2024); Jang et al. (2024); Singh and Jaggi (2020), as a survey paper, the main focus of this paper is to offer a systematic analysis of why and how weight-averaged model-merging  Wortsman et al. (2022) can work from the abovementioned three interesting viewpoints. Based on our viewpoints and findings, we also conclude with useful proposals to provide a good reference for the community.
@@ 2 Related Work
Figure 1: Average of all images for each class (a) and the linear classifier visualization for each class on Cifar100 Krizhevsky (2009) dataset (b). We train a single linear classifier for all classes then wrap it back to the size of input images. The detailed experimental analysis is in Sec. 6.1.
@@
In recent years, model-merging Sung et al. (2023); Sanjeev et al. (2024); Daheim et al. (2023); Navon et al. (2023); Singh and Jaggi (2020); Yang et al. (2023a); Jang et al. (2024); Almakky et al. (2024); Singh et al. (2024) has attracted growing interest. This approach has led to a range of methods focused on merging weights from independently trained models to enhance model performance or generalization. For example, Wortsman et al. Wortsman et al. (2022) proposed an approach of averaging, or greedily averaging, the parameters from various models to yield “Model Soups” that show more accurate performance than the original models used to produce the “soup” but without introducing more computations into inference. Further, Yadav et al. proposed Ties-Merging Yadav et al. (2024) based on parameter averaging by considering conflicts between the model weights, allowing for smooth parameter merging across models. Some other approaches Ainsworth et al. (2022); Entezari et al. (2021); Jordan et al. (2022) address the problem from a different perspective, focusing on taking weight merging operations via finding a good way to interpolate on the loss basins. Zipit Stoica et al. (2023) takes one step ahead from Git Re-Basin Ainsworth et al. (2022) to align weights from inter-model and intra-model based on the alignment of features between merging models. Yu et al. Yu et al. (2024) proposed a DARE model to combine models trained on domain-specific data to create a domain-agnostic model-merging that generalizes well across diverse distributions. This approach ensures that domain knowledge is preserved while enhancing the model’s ability to generalize across different tasks. Beyond model-merging, other techniques, such as mode connectivity path ensembles Garipov et al. (2018), analyze pathways in the parameter space between models. By identifying optimal modes along these paths, these models help improve the performance after merging. Fisher Merging Matena and Raffel (2022) selects parameters that approximately maximize the joint likelihood of the models’ parameter posteriors.
Despite these advances, research into the underlying mechanisms that make model-merging effective remains limited. This gap in understanding serves as the motivation for our paper.
@@ 3.1 Connections Between Inner Product and Template Matching
The weights of the model can be interpreted as templates or prototypes for the respective class activations and it has a deep connection with machine learning. A couple of decades ago, Adaboost Freund et al. (1999) contained heuristically designed patterns such as HAAR features Viola and Jones (2001), which are simple hand-crafted filters that capture edges/lines/other low-level image features through weighted sums and differences in pixel intensities. During training, Adaboost iteratively adjusts the weights of weak classifiers to improve their performance on misclassified examples.
In a linear classification, once the weight matrix is learned, we wrap it back to have the same size as the input images. Take Cifar100 for instance, each image within the dataset is 32$\times$32$\times$3, if we flatten it and it will become a 3072$\times$1 vector; the weight matrix size is 100$\times$3072 without considering bias; in this case, each row of the weight matrix corresponds to each class to be classified and thus each row vector (1$\times$3072) can be wrapped back to 32$\times$32$\times$3 with the same size of the original images. The learned weight matrix serves as simplified templates that represent each class within the data, as shown in Fig. 6.1.
A key mechanism enabling these templates to operate effectively is the use of inner products to apply them to the input data. When an image is presented to the model, its activation is determined by the inner product between the image features and per class weight vector. For each class, which weight vector yields the highest activation is selected as the predicted class of that vector, indicating that the “template” aligns (or similar to) the input image most closely. This phenomenon also holds for deeper networks, but differently from the linear classifier, weight templates in each layer are learned to detect different features that contribute the most towards the objectives (e.g., in the lower layer, low-level features, e.g., edge/line features, are tend to be learned). The averaging operation in the weight space is represented as Fig. 2, which can be seen as the Mixup Zhang (2017) operation in the weight space. Therefore, the linear combinations in the weight space have a similar effect as in the feature space. This also explains why recent research Dekoninck et al. (2023); Zhang et al. (2023); Ilharco et al. (2022); Zhou et al. ; Ortiz-Jimenez et al. (2023) applies “model arithmetic” on multiple models in the weight space can achieve the effect of feature arithmetic as Word2vec Mikolov (2013).
@@ 3.2 Formal Representation from the Perspective of Template Matching
where $\|\mathbf{w}\|$ and $\|\mathbf{x}\|$ are the norms of $\mathbf{w}$ and $\mathbf{x}$, and $\theta$ is the angle between the two vectors. When $\theta=0$, then $\cos\theta=1$, which means that we have $\mathbf{w}\cdot\mathbf{x}=\|\mathbf{w}\|\|\mathbf{x}\|$, resulting in a weight matrix $\mathbf{w}$ that perfectly matches the input $\mathbf{x}$, as the inner product in (1) is maximized when the template $\mathbf{w}$ is perfectly aligned with the input $\mathbf{x}$. On the other hand, when $\theta\neq 0$, then $\cos\theta<1$, which means that the template and image features become misaligned. This mechanism also applies to the layers within deep learning models. The visualization of deep models can be found in the supplementary materials.
@@
Figure 3: (a) The 10 model-merging/ensemble accuracy comparison between different architectures on Cifar100. Comparison between averaging on Weights (Uniform Soups, Greedy Soups), averaging on features (Ens Logits, Ens Features, Greedy Ens Logits, Greedy Ens Features) and original model performance (Performance Average) on Cifar100 using many different deep learning architectures, including CNNs and ViTs. The detailed experimental analysis is in Sec. 6.2. (b) Scatter plot comparing the classification accuracies of model merge (uniform soup) and model ensemble (logits ensemble) over different weights magnitudes with VGG19 model on Cifar100. The values on the X and Y axis represent the accuracies. The experimental analysis is in Sec. 6.3.
In this section, we explore the differences between averaging on weights and features by analyzing their effects on the model effectiveness and efficiency, one example experiment is shown in Fig. 3(a).
@@ 4.2 Non-linear Model Scenarios
In such a case, the equivalency between $\overline{\mathbf{h}}$ and $\overline{\mathbf{h}}^{\prime}$ seen in the linear models may not hold for non-linear models. Take ReLU for instance, if elements of $\mathbf{W}_{1}\mathbf{x}$ and $\mathbf{W}_{2}\mathbf{x}$ exhibit totally opposite activations (same absolute values but with different signs), where their average will tend towards zero, leading to a large loss of information after the non-linear activation. Therefore, averaging on weights tends to smooth out the contributions from $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, which can be beneficial in certain scenarios by reducing variance or providing a regularizing effect. However, this smoothing may also lead to a loss while extracting fine-grained details. In contrast, averaging on features allows each weight matrix to independently influence the final output after activation. However, averaging on features (model ensemble) increases the computational costs linearly with the number of ensembled models, while averaging on weights (model-merging) does not suffer. The FLOPs of models considered in the paper can be found in the supplementary materials.
@@ 5 The Model Predictions and Weight Magnitudes
In this section, we will examine how the weight-averged merging behaves when the candidate model weights are multiplied by different multiples (e.g., 90×, 100×, etc.), defined as magnitude factors. This helps us understand how robust model-merging works.
@@ 5.1 Analysis of Weight Magnitude After Averaging
We start by analyzing the effect of averaging the weights of two independent matrices, denoted as $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, on their respective magnitudes. We define the maximum entry of both matrices as $M=\max(\mathbf{W}_{1},\mathbf{W}_{2})=\max_{i,j}(|(\mathbf{W}_{1})_{i,j}|,|(\mathbf{W}_{2})_{i,j}|)$. When we compute the averaged weights, we obtain:
@@
Mathematical Justification. Using the triangle inequality Tversky and Gati (1982), we establish:
@@
Thus, it follows that $\|\overline{\mathbf{W}}\|\leq M$ combining with Eq. 10. It confirms that averaging the weights results in a new weight whose entries’ upper bound is equal to the maximum entry of the original weights. Thus, the operation of averaging the weights $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ typically results in a reduction of the maximum weight magnitude. This conclusion also extends to complex deep neural networks.
@@ 5.3 How Weight Magnitudes and Variance Effects Model Outputs
We have the following property 1 and theorem 1 to build the connections between model weights’ magnitudes/variances and outputs’ magnitudes/variances, following  Wang et al. (2024):
@@
where $M_{l}$ denotes the maximum number of layers, and $L$ represents the Lipschitz constant. This inequality shows that the output variance grows with the depth of the network, and is impacted by the variances of the weights/bias.
@@
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 3: As the aforementioned analyses and experimental study in Sec. 6.3, model-merging is more robust to the changes in weights’ magnitudes/variances than model ensemble, by bringing down the maximum magnitudes/variances of individual models before merging. This is because, generally, the small magnitudes of weights enhance robustness by minimizing the network’s sensitivity towards input variations, but somehow reduce the expressiveness from ensembling all models. From this perspective, model-merging acts similarly to regularization. This balance between robustness and expressiveness is crucial and should be carefully considered based on the specific task requirements. For certain tasks where noise resilience is more valuable than fine-grained feature sensitivity, lower weight magnitudes can improve performance.
@@ 6.1 The Patterns Contained in Model Weights
Figure 4: Average of all images for each class and linear classifiers visualization for each class in Tiny ImageNet Le and Yang (2015) dataset.
As shown in Fig. 1(a,b), the visualization of the linear classifier on Cifar100 explicitly shows the patterns per classifier learned, which are strongly correlated with the average of images per class. The patterns learned per class act as a fused representative image. Interestingly, if we merge the weights of the first 5 classes with the last 5 classes in fig. 1, it gives us the results of fig. 2. As mentioned, the weight-averaged model-merging acts as a Mixup operation, which is a linear combination of different matrices, in the weight space. For instance, when the “bicycle” weight matrix is combined with the “bottle” weight matrix, we can see the wheels of bikes combined with a vertical bottle. We can also observe that, after merging, each element within the combined template noticeably fades. This phenomenon can also explain the limitation of expressiveness of the merged model as described in section 4.
fig. 4 visualizes the image average per class and its corresponding weight for Tiny ImageNet classes. For instance, “goldfish” shows a reddish center in the averaged images and also shows a reddish center in the classifier weights; similarly, we can observe the eyes and nose of “koala”; for pretzel, we can see the shape of the cracker from its weight visualization. However, they are not as clear as for Cifar100. This may be due to the relatively large dataset size, as the average of all images in a class itself is also vaguer. This compromised clarity can also be caused by the limited expressiveness of single linear classifiers and there are more complicated patterns present in Tiny ImageNet than Cifar100. Besides the linear classifier visualization, we also conduct common deep learning models with nonlinear weights visualization by viewing the first layer of convolutional kernels (7 $\times$ 7 in size) from the ResNet50 model He et al. (2016) in fig. 7 (VGG model weight visualization can be found in the supplementary materials). We can see the patterns of texture and edges for the basic level geometric feature detection is similar to the Haar features described in Sec. 3.1.